OpenStudy (anonymous):

solve: (x-1/3)(x-1/3)+(x-1/3)(x+2/9)=0

OpenStudy (anonymous):

factor first

OpenStudy (anonymous):

3(x-1/3)+(x+2/9)=0

OpenStudy (anonymous):

3x-1+x+2/9=0

OpenStudy (anonymous):

\[(x-\frac{1}{3})(x-\frac{1}{3})+(x-\frac{1}{3})(x+\frac{2}{3})\] \[(x-\frac{1}{3})(x-\frac{1}{3}+x+\frac{2}{3})\] \[(x-\frac{1}{3})(2x+\frac{1}{3})=0\]

OpenStudy (anonymous):

4x-0.7=0

OpenStudy (anonymous):

4x=0.7 x=.175

OpenStudy (anonymous):

\[x=\frac{1}{3}\] or \[2x=-\frac{1}{3}\] \[x=-\frac{1}{6}\]